OpenStudy (anonymous):

Assuming the balls to be identical except for difference in colours, the number of ways in which one or more balls can be selected from 10 white, 9 green and 7 black balls is : (1) 880 (2) 629 (3) 630 (4) 879

OpenStudy (anonymous):

\[(10+1) \times (9+1) \times (7+1) -1 = 879 \]

OpenStudy (anonymous):

can you explain it properly? why not 10 x 9 x 7 ??

OpenStudy (anonymous):

For \(n \) identical balls (balls of same color) you can select none or more them in \( (n+1)\) ways.

OpenStudy (anonymous):

okay...then why -1 in the end?

OpenStudy (anonymous):

Because this include the case where we select none from all of the groups but the question ask for one or more, hence we need to subtract 1 at last.

OpenStudy (anonymous):

Ah! I see. Thanks a lot! :)